Prepare 100 mL of 0.200 bicarbonate buffer, pH 10 by calculating the mass of sodium bicarbonate and sodium carbonate.

    a. Use the Henderson-Hasselbalch equation to calculate the volume of each stock

        solution needed.  pH = pKa + log [conjugate base] / [weak acid]1

    b. Check your calculations with other students. See the instructor if there is uncertainty.

    c. Make the solution and check the pH of a portion of your buffer solution using the pH meter.

                                              L = 245μL 



1
Expert's answer
2021-09-30T03:04:58-0400

100 ml × 0.2 = 20 ml


Molar mass of sodium carbonate = 105.9888 g/mol


Molar mass of sodium bicarbonate = 84.007 g/mol


= 105.9888 × 84.007/20


= 4.452 × 102


= log 4.452 × 102 = 2.65


pH = 10 + 2.65 = 12.65